OpenStudy (anonymous):

Find the exact value of the expression given that secx=3/2, cscy=3 and x and y are in quadrant 1

OpenStudy (anonymous):

Someone help!

OpenStudy (anonymous):

CALLISTO HELP!

OpenStudy (callisto):

If secx=3/2, what is cosx?

OpenStudy (anonymous):

2/3??

OpenStudy (callisto):

Yes. When cos x = 2/3, what is x?

OpenStudy (anonymous):

I dont know??

OpenStudy (callisto):

Take inverse of cosine for both sides.

OpenStudy (anonymous):

can u please write this out i am very confused!

OpenStudy (callisto):

Inverse of cosine => \(\cos^{-1}\)

OpenStudy (anonymous):